OpenStudy (anonymous):

Question: A smartphone currently is worth $175 and depreciates at 25% per year. What will the value be in 3 years? What equation do I use to solve this?

OpenStudy (anonymous):

loses 25% of its value means it retains 75% of its value use \(175\times .75^t\)

OpenStudy (anonymous):

eqquation close to p=(1+r/100)^n n=number of years compound interest

OpenStudy (anonymous):

or if you prefer \(175\times (\frac{3}{4})^t\) either way you want \(175\times (.75)^3\) and a calculator is needed for this one

OpenStudy (anonymous):

well done satellite

OpenStudy (anonymous):

Thanks guys.(: